log x = log6 36 + log81 9 – log2 4

Answer by nerdybill(7384)

log x = log6 36 + log81 9 – log2 4
we can rewrite as:
log x = log6 6^2 + log81 81^(1/2) – log2 2^2
Now we simplify:
log x = 2 + (1/2) – 2
log x = 1/2
x = 10^(1/2)
x = 3.162